# The Sparse Frontier: Sparse Attention Trade-offs in Transformer LLMs

### v1

@@ Appendix D FLOPS breakdown
$\text{FLOPS}_{\text{attention}}$ $=N\cdot(2Ld^{2}+2L^{2}d\cdot\rho+3hL^{2}\cdot\rho+2L^{2}d\cdot\rho)$ (5)
$\text{FLOPS}_{\text{mlp}}$ $=N\cdot(4Ld\cdot d_{\text{mlp}}+2L\cdot d_{\text{mlp}})$ (6)
@@
Where $L$ is sequence length, $d$ is hidden dimension, $h$ is number of heads, $N$ is number of layers, $d_{\text{mlp}}$ is MLP intermediate dimension, $|V|$ is vocabulary size, and $\rho$ represents attention density (1-sparsity). For decoding the typical generation length (in our case 200 tokens on average) is orders of magnitude shorter than prefilling lengths. Therefore, we estimate the decoding FLOPS by calculating the cost of processing one new token while attending to the entire context of length $L$.
For sparse attention methods, we also account for the computational cost of importance estimation—the process of determining which subset of the attention matrix to compute. For Vertical-Slash pattern, this includes dot products between keys and a subset of queries, softmax operations, aggregation over queries and diagonals, top-k selection, and index building. The indexing cost is approximately:
$\text{FLOPS}_{\text{VS indexing}}$ $=N\cdot h\cdot[2dLq+3Lq+2Lq+2L\log_{2}(L)+\frac{L}{64}(k_{v}+k_{s})]$ (8)
@@
Where $q$ is the number of last queries used for importance estimation, and $k_{v}$ and $k_{s}$ are the number of vertical and slash patterns selected.
For Quest during decoding, the indexing cost involves computing importance scores for each page in the KV cache:
$\text{FLOPS}_{\text{Quest indexing}}$ $=N\cdot h\cdot[2d\cdot\frac{L}{p}+3d\cdot\frac{L}{p}+\frac{L}{p}\log_{2}(\frac{L}{p})]$ (9)
Where $p$ is the page size (16 in our case). These indexing costs are included in our isoFLOPS analysis to ensure fair comparison between dense and sparse attention methods.
Figure 12 illustrates the FLOPS required for both prefilling and decoding operations, revealing two distinct computational regimes in transformer models: a linear-dominated regime at shorter sequences (below 16K) where non-attention components (embedding, MLP, and output layers) determine the computational cost, and a quadratic-dominated regime at longer sequences (beyond 16K-48K, depending on model size) where attention computation with its $O(L^{2})$ complexity becomes the primary factor driving the total computational requirements.
A critical insight emerges in the quadratic-dominated regime: at sufficient sequence lengths, larger models with sparse attention can require fewer FLOPS than smaller dense models. This efficiency crossover occurs because sparsity substantially reduces the attention cost which scales faster than the linear components. This effect becomes more pronounced as sequence length increases, creating an increasingly favorable computational trade-off for sparse large models over dense small ones.
@@ Appendix F Example Story Narrative
Chapter 1:
Beneath gentle breezes, Arion ventured into Athens, curious about its secrets. Long journeys had led Arion to Athens, a step closer to understanding. Soon enough, a tense negotiation seized everyone’s attention. Cleo appeared as if expecting Arion, engaging them without delay. Carefully, they navigated the topic of old feuds, wary of awakening dormant animosities that still simmered. In a calm moment, they compared notes on the traders who passed through Athens, each leaving their subtle mark. In hushed tones, they spoke of local customs and distant rumors, sharing hints of hidden pathways. Following subtle bargaining with Cleo, Arion claimed ownership of lavish crystal lamp. With a light gesture, Arion acknowledged Cleo once more before departing. Nothing would be the same as Arion left Athens, thoughts turning inward. In quiet corners, ambitions simmered, waiting for a spark.
Chapter 2:
At dawn, Arion reached the gates of Hippo Regius, where merchants and travelers converged. This place might hold a clue Arion had long sought. Hardly had Arion arrived before a violent storm stirred uneasy whispers. Thanos approached Arion, eyes bright with opportunity. They lingered over tales of old alliances and forgotten disputes, weaving past into present. They debated the meaning of recent events, each seeking patterns in the chaos. Their reflections turned to the interplay of supply and demand, seeing how fortunes might turn in an instant. After reaching terms with Thanos, Arion took possession of ceremonial gold seal. Arion turned from Thanos, ready to move on. In parting, Arion acknowledged that the journey still had far to run. Hidden corners of the city promised knowledge or peril.
Chapter 3:
The threshold of Emerita Augusta welcomed Arion, who felt the weight of untold stories. Arion came here hoping to learn something new, or perhaps gain an advantage. Within hours, a violent storm disrupted the familiar routines. There, Arion encountered Niko, who seemed eager to exchange words or goods. Their words lingered on rumors of distant lands, where fortunes or ruin awaited bold seekers. They debated the meaning of recent events, each seeking patterns in the chaos. Their dialogue danced around subtle clues, each suggestion hinting at treasures undiscovered. The transaction concluded with Arion acquiring delicate porcelain sword from Niko. With a light gesture, Arion acknowledged Niko once more before departing. Eventually, Arion moved on, carrying new impressions forward. The distant hum of voices hinted at unseen deals.
Chapter 4:
Under fading daylight, Arion set foot in Berenice, eager to learn what it offered. A quiet determination brought Arion to Berenice, ever searching for meaning. a sudden market crash cast its shadow over Berenice, changing plans and minds. Roxana approached Arion, eyes bright with opportunity. Together, they reflected on the nature of trust and deceit, aware that fate often twists. They compared accounts of strange visitors bearing knowledge or confusion, each arrival a new riddle in Berenice. A short exchange revealed uncharted corners of Berenice, where knowledge or secrets might dwell. mystic bronze lamp changed hands as Arion completed the purchase from Roxana. Arion handed over lavish crystal lamp to Roxana as the deal closed. With a light gesture, Arion acknowledged Roxana once more before departing. As Arion prepared to depart, the path ahead remained uncertain but compelling. Somewhere, a whisper promised answers for those who dared.
Chapter 5:
Under fading daylight, Arion set foot in Syracuse, eager to learn what it offered. In pursuit of truth, Arion looked to Syracuse for subtle revelations. Not long after arriving, an opulent banquet shook the local order. Phaedra appeared as if expecting Arion, engaging them without delay. Their words traced over delicate negotiations that had once sealed lasting truces in Syracuse. Carefully, they navigated the topic of old feuds, wary of awakening dormant animosities that still simmered. They delved into the subtle art of earning trust in a place where trust was scarce and hard-won. With measured consideration, Arion purchased engraved emerald goblet from Phaedra, examining it closely. In quiet understanding, Arion left Phaedra, their paths diverging. In parting, Arion acknowledged that the journey still had far to run. A subtle tension lingered, as though fate held its breath.

### v2

@@ B.1 Cost Formulas
$\text{FLOPS}_{\text{attention}}$ $=N\cdot(2Ld\cdot(d+2d_{h}n_{kv}+d)+\rho\cdot(2hL^{2}d_{h}+3hL^{2}+2hL^{2}d_{h}))$ (3)
$\text{FLOPS}_{\text{mlp}}$ $=N\cdot(6\cdot L\cdot d\cdot d_{\text{mlp}}+2\cdot L\cdot d_{\text{mlp}})$ (4)
@@
For decoding, which is memory-bound, we measure memory accesses:
$\text{Memory}_{\text{decode}}$ $=\text{Memory}_{\text{weights}}+B\cdot\text{Memory}_{\text{kv\_cache}}$ (6)
$\text{Memory}_{\text{kv\_cache}}$ $=N\cdot 2\cdot L\cdot d_{h}\cdot n_{kv}\cdot\rho$ (7)
$\text{Memory}_{\text{weights}}$ $=N\cdot(4d^{2}+3d\cdot d_{\text{mlp}})+d\cdot|V|+d$ (8)
where $L$ is sequence length, $d$ is hidden dimension, $h$ is number of query heads, $d_{h}$ is head dimension, $n_{kv}$ is number of key-value heads, $N$ is number of layers, $d_{\text{mlp}}$ is MLP intermediate dimension, $|V|$ is vocabulary size, $\rho$ represents attention density ($1-\text{sparsity}$), and $B$ is batch size.
For sparse methods, we include importance estimation overhead. Vertical-Slash indexing:
$\text{FLOPS}_{\text{VS indexing}}$ $=B\cdot N\cdot h\cdot[2dLq+3Lq+2Lq+2L\log_{2}(L)+\frac{L}{64}(k_{v}+k_{s})]$ (9)
Quest indexing requires loading page representations:
$\text{Memory}_{\text{Quest indexing}}$ $=B\cdot N\cdot n_{kv}\cdot 2d\cdot\frac{L}{p}$ (10)
@@
where $q$ is the number of queries for importance estimation, $k_{v}$$k_{s}$ are selected vertical/slash patterns, and $p$ is page size (16).
B.2 Prefilling: Sequence Length Drives Sparsity Impact
[…]
Attention cost scales quadratically with sequence length ($O(L^{2})$) during prefilling, while non-attention costs (MLP, embeddings, logits) scale linearly ($O(L)$). This creates two regimes: at shorter sequences, non-attention components dominate; at longer sequences, attention becomes the primary cost.
[…]
B.3 Decoding: Sequence Length and Batch Size Both Matter
[…]
Unlike prefilling, decoding cost depends on both sequence length and batch size. KV cache access scales linearly with context length ($O(L)$) and batch size ($O(B)$), while model weight loading is constant. This creates an important distinction: for prefilling, all cost components scale linearly with batch size, so the attention-to-total ratio remains constant. Decoding behaves differently—model weights are loaded once per forward pass regardless of batch size, while KV cache access scales with batch size.
[…]
B.4 Sliding-Window Architectures Need Longer Sequences
Figure 14: Attention cost ratio comparison between Qwen 14B (dense attention) and Gemma 12B (sliding-window attention) at batch size 8. Gemma’s architectural sparsity results in substantially lower attention ratios, requiring longer sequences for additional sparse attention to provide meaningful cost reduction.
Models with built-in architectural sparsity, such as Gemma 3’s sliding-window attention (5 out of 6 layers use 1024-token windows), have lower baseline attention ratios. Figure 14 compares similar-sized models: at 64K tokens with batch size 8, Qwen 14B has 76% attention ratio for prefilling versus Gemma 12B’s 42%. For decoding, Qwen reaches 79% versus Gemma’s 61%. At 128K tokens, Gemma’s attention ratio rises to 54% for prefilling and 75% for decoding.
This difference has practical implications for isoCost comparisons. Because Gemma has a lower attention-to-total ratio at 64K, sparse prefilling reduces a smaller fraction of total FLOPs than for Qwen, so the dense–sparse frontiers overlap later (i.e., at higher costs / longer sequences). In principle, one would expect stronger overlap for Gemma at longer contexts as the attention ratio rises, but in our experiments most model–task configurations at 128K exhibit near-zero performance, preventing a meaningful prefilling comparison at that length. Decoding is less constrained: at sufficiently high batch size (e.g., $B=64$), KV-cache transfers dominate even for sliding-window models, so sparse decoding still yields practical benefits and can exhibit dense–sparse overlap in isoCost space (see Section 4).
Appendix C Comparison to Prior Work
Several concurrent works have explored evaluation of sparse attention methods. We summarise the key differences below.
SCBench (Li et al., 2025)
does not control for sequence length and evaluates at most two models from a single family, making it difficult to analyse the effects of sequence length and model size on sparse attention performance. Our work systematically varies sequence length (16K–128K tokens) and evaluates multiple model sizes within each of three model families.
Liu et al. (2025b)
only considers models up to 10B parameters and does not address sparse attention in the prefilling phase. In contrast, we evaluate models up to 72B parameters and analyse both prefilling and decoding phases separately, revealing phase-specific behaviours.
Yuan et al. (2024)
tests sequence lengths only up to 32K tokens and includes models up to 10B parameters. Our evaluation extends to 128K tokens and 72B parameters, capturing the regime where sparse attention benefits are most pronounced.
In summary, our work is the first to systematically conduct an isoCost analysis for sparse attention, providing new insights into efficiency–accuracy trade-offs and generalisation across model size, sequence length, and sparsity.
Appendix D Extra Results
D.1 Statistical Error Bounds
We report standard error in Figure 2 (main text, RQ2) for our per-task and per-method results. We omit standard error bars in isocost figures (Figure 1) for visual clarity, as the standard error is negligible. Here we derive the upper bound for the standard error across all configurations.
Since performance metrics lie in the $[0,1]$ range, the maximum standard deviation is $\sigma_{\max}=0.5$ (achieved when the metric has a Bernoulli distribution with $p=0.5$). For configurations where we aggregate results over $N$ samples, the standard error is:
$SE=\frac{\sigma}{\sqrt{N}}\leq\frac{\sigma_{\max}}{\sqrt{N}}=\frac{0.5}{\sqrt{N}}$ (11)
In Figure 1, we aggregate performance across 9 tasks with 100 samples each (for Qwen), yielding $N=900$ samples total:
$SE_{\max}=\frac{0.5}{\sqrt{900}}=\frac{0.5}{30}\approx 0.0167$ (12)
This upper bound of approximately $0.017$ is substantially smaller than the performance differences we observe between configurations (typically >0.05>0.05), justifying our decision to omit error bars for visual clarity in the isocost analysis.
D.2 Per-Task Results by Model Family
This section provides per-task performance breakdowns for each model family, complementing the aggregated analysis in Section 4.2. Figures 15, 16 and 17 show results for Qwen 2.5, Llama 3.1, and Gemma 3 respectively.
Figure 15: Per-task performance for Qwen 2.5 models (7B, 14B, 32B, 72B) at sequence lengths 16k, 32k, and 64k. Left: prefilling methods. Right: decoding methods.
Figure 16: Per-task performance for Llama 3.1 models (8B, 70B) at sequence lengths 16k, 32k, and 64k. Left: prefilling methods. Right: decoding methods.
Figure 17: Per-task performance for Gemma 3 models (4B, 12B, 27B) at sequence lengths 16k, 32k, and 64k. Left: prefilling methods. Right: decoding methods.
D.3 Sequence Length Effects
Figure 18 presents the absolute error perspective on sequence length effects, complementing the relative error analysis in Section 4.3. The absolute error is $\bar{p}_{\text{dense}}-\bar{p}_{\text{sparse}}$, where $\bar{p}$ denotes mean performance. The pattern mirrors the relative error findings: longer sequences tolerate higher sparsity with smaller absolute performance degradation.
Figure 18: Absolute error vs. sparsity across sequence lengths. Results aggregated across all tasks, methods, and models (Qwen 2.5, Llama 3.1, Gemma 3).
Figure 19 provides per-family breakdowns of the sequence length analysis. The trend of improved sparsity tolerance at longer sequences holds consistently across all three model families, with minor variations in magnitude.
Figure 19: Sequence length effects on sparsity tolerance by model family. Top row: absolute error vs. sparsity. Bottom row: relative error vs. sparsity. Results aggregated across all tasks and methods within each family.
D.4 Model Size Analysis
We analyse how sparsity tolerance varies with model scale. Figure 20 shows model size effects aggregated across all tasks, methods, and sequence lengths for each model family. On average, model size shows no clear correlation with sparsity tolerance—the lines for different model sizes largely overlap, indicating that larger models do not systematically tolerate more or less sparsity than smaller ones.
Figure 20: Model size effects on sparsity tolerance aggregated across all tasks. Top row: absolute error vs. sparsity. Bottom row: relative error vs. sparsity. Results aggregated across all tasks, methods, and sequence lengths 16–64k for each model family.
However, this aggregate finding masks important task-dependent patterns revealed in Figure 21.
Figure 21: Model size effects on sparsity tolerance for Qwen 2.5 (7B–72B). Absolute error is $\bar{p}_{\text{dense}}-\bar{p}_{\text{sparse}}$; relative error is $(\bar{p}_{\text{dense}}-\bar{p}_{\text{sparse}})/\bar{p}_{\text{dense}}$, where $\bar{p}$ denotes mean performance. Top row: absolute error vs. sparsity. Bottom row: relative error vs. sparsity. Left column: easy tasks (Story Retrieval, Ruler NIAH). Right column: hard tasks (Ruler VT, Story Filtering). Results aggregated across methods and sequence lengths 16–64k.
Model size effects depend on task difficulty.
Figure 21 presents contrasting perspectives on model size effects. On tasks where all model sizes achieve near-perfect dense accuracy (left column: Story Retrieval, Ruler NIAH), larger models tolerate more sparsity—at sparsity 0.95 ($1/20$ budget), 72B shows 0.20 absolute error compared to 0.50 for 7B. Conversely, on challenging tasks where dense accuracy scales with model size (right column: Ruler VT, Story Filtering), larger models exhibit larger absolute errors at equivalent sparsity—72B shows 0.21 absolute error while 7B shows only 0.03 at sparsity 0.95. The relative error perspective (bottom row) shows consistent patterns: larger models have lower relative error on easy tasks but higher relative error on hard tasks.
These divergent patterns arise from how sparsity interacts with model capacity. Sparse attention reduces effective model capacity by limiting information flow. When a model operates far above a task’s difficulty threshold, this capacity reduction has minimal impact on outputs. When model capacity approximately matches task difficulty, even modest sparsity degrades performance. Larger models achieve higher dense accuracy on difficult tasks, operating closer to their capacity limits on these tasks—making them more vulnerable to capacity reductions from sparsity. Evaluations on tasks where models achieve perfect or near-perfect accuracy—common in benchmarks like Needle-in-a-Haystack—cannot reveal these vulnerabilities.
@@ Appendix F Example Story Narrative
Chapter 1: Beneath gentle breezes, Arion ventured into Athens, curious about its secrets. Long journeys had led Arion to Athens, a step closer to understanding. Soon enough, a tense negotiation seized everyone’s attention. Cleo appeared as if expecting Arion, engaging them without delay. Carefully, they navigated the topic of old feuds, wary of awakening dormant animosities that still simmered. In a calm moment, they compared notes on the traders who passed through Athens, each leaving their subtle mark. In hushed tones, they spoke of local customs and distant rumors, sharing hints of hidden pathways. Following subtle bargaining with Cleo, Arion claimed ownership of lavish crystal lamp. With a light gesture, Arion acknowledged Cleo once more before departing. Nothing would be the same as Arion left Athens, thoughts turning inward. In quiet corners, ambitions simmered, waiting for a spark. Chapter 2: At dawn, Arion reached the gates of Hippo Regius, where merchants and travelers converged. This place might hold a clue Arion had long sought. Hardly had Arion arrived before a violent storm stirred uneasy whispers. Thanos approached Arion, eyes bright with opportunity. They lingered over tales of old alliances and forgotten disputes, weaving past into present. They debated the meaning of recent events, each seeking patterns in the chaos. Their reflections turned to the interplay of supply and demand, seeing how fortunes might turn in an instant. After reaching terms with Thanos, Arion took possession of ceremonial gold seal. Arion turned from Thanos, ready to move on. In parting, Arion acknowledged that the journey still had far to run. Hidden corners of the city promised knowledge or peril. Chapter 3: The threshold of Emerita Augusta welcomed Arion, who felt the weight of untold stories. Arion came here hoping to learn something new, or perhaps gain an advantage. Within hours, a violent storm disrupted the familiar routines. There, Arion encountered Niko, who seemed eager to exchange words or goods. Their words lingered on rumors of distant lands, where fortunes or ruin awaited bold seekers. They debated the meaning of recent events, each seeking patterns in the chaos. Their dialogue danced around subtle clues, each suggestion hinting at treasures undiscovered. The transaction concluded with Arion acquiring delicate porcelain sword from Niko. With a light gesture, Arion acknowledged Niko once more before departing. Eventually, Arion moved on, carrying new impressions forward. The distant hum of voices hinted at unseen deals. Chapter 4: Under fading daylight, Arion set foot in Berenice, eager to learn what it offered. A quiet determination brought Arion to Berenice, ever searching for meaning. a sudden market crash cast its shadow over Berenice, changing plans and minds. Roxana approached Arion, eyes bright with opportunity. Together, they reflected on the nature of trust and deceit, aware that fate often twists. They compared accounts of strange visitors bearing knowledge or confusion, each arrival a new riddle in Berenice. A short exchange revealed uncharted corners of Berenice, where knowledge or secrets might dwell. mystic bronze lamp changed hands as Arion completed the purchase from Roxana. Arion handed over lavish crystal lamp to Roxana as the deal closed. With a light gesture, Arion acknowledged Roxana once more before departing. As Arion prepared to depart, the path ahead remained uncertain but compelling. Somewhere, a whisper promised answers for those who dared. Chapter 5: Under fading daylight, Arion set foot in Syracuse, eager to learn what it offered. In pursuit of truth, Arion looked to Syracuse for subtle revelations. Not long after arriving, an opulent banquet shook the local order. Phaedra appeared as if expecting Arion, engaging them without delay. Their words traced over delicate negotiations that had once sealed lasting truces in Syracuse. Carefully, they navigated the topic of old feuds, wary of awakening dormant animosities that still simmered. They delved into the subtle art of earning trust in a place where trust was scarce and hard-won. With measured consideration, Arion purchased engraved emerald goblet from Phaedra, examining it closely. In quiet understanding, Arion left Phaedra, their paths diverging. In parting, Arion acknowledged that the journey still had far to run. A subtle tension lingered, as though fate held its breath.
@@ 2. Then, provide your final answer following this exact format:
Appendix H Use of AI Assistants
We used Claude Opus 4.5 for grammar and style suggestions during the writing of this paper. All scientific content, analysis, and conclusions are the authors’ own work.